OpenStudy (anonymous):

Visible light has a wavelength of about 0.5 microns. What is its frequency in Hz (cycles per second)?

OpenStudy (anonymous):

0.5 microns = 5.0 × 10^-7 meters C = 300,000,000 m/s = frequency * wavelength frequency = 300,000,000/5.0 × 10^-7 frequency = 6.0 x 10^14 Hz